OpenStudy (anonymous):

Find the domain of f for each of the following functions. If you need to enter - or , enter -INFINITY or INFINITY: f(x)=sqrt(4x+8)

OpenStudy (anonymous):

\[4x+8\geq0\] \[4x\geq-8\] \[x\geq -2\]

OpenStudy (anonymous):

interval notation is \[[-2,\infty)\]

OpenStudy (anonymous):

THANK YOU SO MUCH (: